OpenStudy (anonymous):

Find the equation of a straight line that has a slope of and is passing through the point . Simplify and write the equation in slope intercept form

OpenStudy (anonymous):

a line equation you could follow is \[y _{1}- y = m (x _{1} - x)\]

OpenStudy (anonymous):

oops i got that wrong.

OpenStudy (anonymous):

i mean \[y - y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

so m is the slope. y1 is the y coordinate of the given point. and x1 is the x coordinate of the given point.

OpenStudy (anonymous):

you just plug in the numbers and solve for y.

OpenStudy (anonymous):

simply and write the equation in slope intercept form

OpenStudy (anonymous):

so if you plug in the numbers and solve for y, y should equal mx + b

OpenStudy (anonymous):

i don't know how to write it in simple form

OpenStudy (anonymous):

i think the slope intercept form, y=mx + b, is the already the simplified form (:

OpenStudy (anonymous):

how would i write it m=2/3 -2,-4 in slope intercept form

OpenStudy (anonymous):

if you plug in those numbers to the first equation i gave you, you would get y+4=(2/3)(x+2)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

to simplify that, you multiply the (2/3) with x and with 2. then you subtract 4 from both sides. no problem !

OpenStudy (anonymous):

what

OpenStudy (anonymous):

if you want to get y by itself on the left side. thats what you would do. so y= (2/3)x - (8/3) that is your final answer.